No additional news from Production team. A few production jobs are submitted and skim/merge jobs only processed.
Not mentioned about skimming. Assume situation is the same as reported in 8/12.
The location of the samples at KEKCC and on the grid are detailed on the Phase 3 page: https://confluence.desy.de/display/BI/Phase+3+data. The samples are separated into proc9, which includes everything up to exp 8, run 1835 (including the off-resonance sample), and bucket7, which includes exp 8 runs 1836-3123.
The next reprocessing, proc10, will use release-04 (pending release next month) and will feature significant updates for some calibration code.
Beam backgrounds and validation samples:
New beam background overlay files for release-04 have been prepared. Details are available at https://confluence.desy.de/display/BI/Beam+background+samples.
The DP group is producing validation samples for release-04. Details are given in https://agira.desy.de/browse/BII-5150 and the associated sub-tickets.
New, run-dependent MC samples, approximately corresponding to proc9, have been produced. Details are available on https://confluence.desy.de/display/BI/Data+Production+Run+Dependent+MC. It is important to note that beam background overlay samples are only available for a subset of runs and only for the pseudorandom trigger (not delayed Bhabhas). It is also the case that not all detectors are ready for run-dependent MC (a goal for release-04). A total of approximately 10/fb equivalent generic samples will be produced, along with signal sample upon request.
Most MC12 samples are now finished. Additional signal requests will be accepted, but jobs are not likely to saturate the system for a while, at least.
Therefore, only limited number of sites are expected to run jobs
Analysis skimming (uDSTs)
The first skim campaign for MC12 is now complete. Details of available analysis skims are given at https://confluence.desy.de/display/BI/MC12+Skim+Production. The second MC12 skim campaign, which will include charm and semileptonic FEI skims, is currently under preparation and will be processed soon. The Phase III data will also be skimmed, as soon as proc9 is ready on the grid.
2018-07-02 No Donetransfer, several scheduled and rapid increase of Waiting replication
GGUS ticket https://ggus.eu/?mode=ticket_info&ticket_id=142410 has submitted
SE Health check by DDM : download does not work since 2019-05-16 07:11:21 UTC.
Cotinuous timeout failure between NTU-CC-TMP-SE and CNAF-TMP-SE
SE Health check by DDM : download, upload do not work since 2019-09-28 10:53:28 UTC.
SE Health check by DDM
GGUS ticket https://uggus.eu/index.php?mode=ticket_info&ticket_id=137825 has submitted 2018-10-18 14:21
No new assignment of MC production data blocks to this destination
Date, Issue, Tickets...
transfer rate to be zero
McGill-TMP-SE will be decomissioned in early 2018.
Date, Issue, Tickets...
File transfer failures to Napoli-TMP-SE and KEK2-TMP-SE, BNL-TMP-SE observed from 2018-12-05 02:00
GGUS ticket https://ggus.eu/index.php?mode=ticket_info&ticket_id=138882
File Transfer failures : File Transfer Efficiency is too low from UVic-DATA-SE. since about 2018-12-18 1:00 (UTC)
This is a test site. Do not need to report any issue.
Banned as currently no resource behind the CE
Health checker info. : "Short pilot jobs" has been found at 22:20:00 UTC on 2019/09/24.
Job status check: Application finished with errors (5% of the jobs) at 11:15 UTC on 2018/12/21.
"Failed to install DIRAC on " has been found since 20:20:00 UTC on 2018/11/03.
AID: "Aborted Pilot" has been observed since 2019-10-16 22:38 UTC (for 16 hours): JIRA ticket created
Health checker info. : "Short pilot jobs" has been found since 21:20:00 UTC on 2019/06/05.(details)
Health checker info. : "Short pilot jobs" has been found since 04:20:00 UTC on 2019/10/03.
Site is currently Banned due to hardware problem since 2019-07-05
Health checker info. : "BLAH ERROR" has been found since 06:20:00 UTC on 2018/10/19.(details)
Job submission check : Pilot submission failure has been found since 21:25:00 UTC on 2019/02/01.